# The Cone Restriction: An old Approach Revisited

### v5

@@ 2.5 Transverse Equidistribution Estimates
For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, $V^{+}$ is defined by
@@
Now we cover $B_{\rho}$ by finitely overlapping $r^{1/2+{\delta}_{m}}$-ball $B$. For some suitable constant $c>0$, if $cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{\neq}}\emptyset$, then we choose some $x\in cB\cap B_{\rho}\cap\boldsymbol{\mathrm{Z}}$, and use $V_{B}$ to denote $T_{x}\boldsymbol{\mathrm{Z}}$. Define
@@
${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\boldsymbol{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
From the definition above, if $B\in{\mathcal{B}}_{b}$, then $\angle\left(M\left(\theta\right),V^{\bot}_{B}\right)$ may be small for some $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$, and thus the transverse equidistribution estimates may not hold for a function concentrated on ${\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$. Fortunately, there are only a few $\theta$ such that $\left(\theta,v,l\right)\in{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}$ for some $v,l$.
@@
$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\boldsymbol{\mathrm{Z}},B}\neq\emptyset\right\}$
@@
Suppose $A$ is large enough. By the definition of $k$-board norm and Lemma 2.25,
${\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{b}}{{\left\|Eg_{\mathrm{tail}}\right\|}^{p}_{BL^{p}_{k,A/2}\left(B\right)}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}\lesssim\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{p}_{2}$ (2.11)
By Lemma 2.18 and (2.11),
${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+{\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.12)
[…]
Transverse Equidistribution Estimates for Essential Part As shown in [15], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$. The precise statement of the transverse equidistribution estimates for cone is:
@@
Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
@@
Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\boldsymbol{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\boldsymbol{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$. For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\boldsymbol{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
@@
here $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
@@
for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={\widetilde{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}$.
@@
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\widetilde{\left(h_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.14)
@@
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\widetilde{\left(g_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}={\left\|{\widetilde{\left(h_{\mathrm{ess,}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right)}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.15)
Combine (2.13), (2.14), and (2.15), one has
@@
2.6 Polynomial Partitioning
[…]
${\left\|F\right\|}_{L^{1}\left(O_{1}\right)}={\left\|F\right\|}_{L^{1}\left(O_{2}\right)},\ \forall O_{1},O_{2}\in\mathrm{cell}\left(P\right)$ (2.16)
See [5] for details.
After a small perturbation and using Sard’s theorem [6], $P$ is guaranteed to be non-singular by relaxing (2.16). The precise statement of the result is as follows.
Theorem 2.28 ([6], Theorem 5.5).
Suppose $F\in L^{1}$ and $d$ is a positive integer. Then, there exists a non-singular polynomial $P$ of degree $O\left(d\right)$ such that $\mathrm{cell}\left(P\right)\sim d^{n}$ and
${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-n}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)},\ \forall O\in\mathrm{cell}\left(P\right)$
Suppose we have a polynomial $P$ described in Theorem 2.28. Let ${\delta}^{\prime}>0$ be a small number. Let $\mathcal{O}$ denote the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)$. By pigeonholing, we see that at least one of two cases hold:
(1).The cellular term dominates, i.e.,
$\sum_{O\in\mathcal{O}}{{\left\|F\right\|}_{L^{1}\left(O\right)}}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
(2).The algebraic term dominates, i.e.,
${\left\|F\right\|}_{L^{1}\left(N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)\right)}\gtrsim{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
[…]
Partitioning over lower dimensional sets Theorem 2.28 is not sufficient to prove Theorem 2.23. We will use a more involved partitioning result, which partitions a general variety in ${\mathbb{R}}^{n}$.
@@ 2.6 Polynomial Partitioning
Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$. Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}$, where $\boldsymbol{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$. Then, at least one of the following cases holds:
@@
Algebraic case. The exists a $\left(m-1\right)$-dimensional transverse complete intersection $\boldsymbol{\mathrm{Y}}\subseteq\boldsymbol{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
Now we apply Theorem 2.29 to the $k$-board norm estimates:
(1). If the cellular case holds, then it follows that
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim d^{m}{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(O\right)},\ \ \forall\ O\in\mathcal{O}$
where $\mathcal{O}$ is the collection of cells produced by Theorem 2.29.
(2). If the algebraic case holds, then it follows that
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right)}\lesssim{\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Y}}\right)}$
where $\boldsymbol{\mathrm{Y}}$ the a $\left(m-1\right)$-dimensional transverse complete intersection produced by Theorem 2.29.
From the fundamental theorem of algebra, one deduces that
@@
3 Finding Polynomial Structure
In this section, we reformulate the induction arguments in [15] as a recursive process by combining the ideas from [11, 12]. The argument contains two different algorithms. In [alg 1] we apply the iterated polynomial partition until the tangent term dominates. This algorithm passes from an $m$-dimensional to an $\left(m-1\right)$-dimensional situation. This algorithm corresponds to the induction on the scale. [alg 2] is established by repeated application of [alg 1]. The iteration ends when we arrive in the minimal dimensional case. Corollary 2.17, the nested polynomial Wolff axioms, is used to establish the bound. This algorithm corresponds to the induction on the dimension.
Without loss of generality, we assume that the error terms, $\mathrm{RapDec}\left(R\right){\left\|.\right\|}_{2}$, is always negligible in every iteration.
3.1 The first algorithm
This algorithm is processed by iterated polynomial partition. At each step of the iteration, one of the 3 cases will happen: cellular case, transverse case and tangent case. If the tangent case happens, or the scale is tiny enough, then the iteration will end.
Input. [alg 1] will take the following things as the input:
$\bullet$ A grain $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$ of dimension $m$.
$\bullet$ A function $f\in L^{\infty}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ which is tangent to $\left(\boldsymbol{\mathrm{Z}},B\left(y,r\right)\right)$.
$\bullet$ An addmissible large integer $A\in\mathbb{N}$.
$\bullet$ In the initial step, we define
$O_{0}:=\left\{B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\boldsymbol{\mathrm{Z}}\right\},\ \ f_{O_{0}}:=f,\ \ {\hslash}_{0}:=\emptyset,\ \ {\rho}_{0}:=r$
Output. At the $j$-th interation, [alg 1] outputs:
$\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning. Let $\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and $\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
$\bullet$ Spatial scale ${\rho}_{j}$: The ${\rho}_{j}$ is determined by the initial scale $r$ and ${\hslash}_{j}$. Define an auxiliary parameter ${\widetilde{\delta}}_{m-1}$ by
$\left(1-{\widetilde{\delta}}_{m-1}\right)\left(\frac{1}{2}+{\delta}_{m-1}\right)=\frac{1}{2}+{\delta}_{m}$
One deduces that
$\delta_{m-1}/2\leq{\widetilde{\delta}}_{m-1}\leq 2{\delta}_{m-1}$
$\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$. Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)\$.
$\bullet$ A collection of functions ${\left(f_{O_{j}}\right)}_{O_{j}\in{\mathcal{O}}_{j}}$. For each cell $O_{j}$ there exists a translate ${\boldsymbol{\mathrm{Z}}}_{O_{j}}\boldsymbol{:}=\boldsymbol{\mathrm{Z}}\boldsymbol{\mathrm{+}}x_{{\mathcal{O}}_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $\left({\boldsymbol{\mathrm{Z}}}_{O_{j}},B_{O_{j}}\right)$
$\bullet$ A large integer $d$ which depends only on the admissible parameters and $\mathrm{deg}\boldsymbol{\mathrm{Z}}\$}.
$C^{\mathrm{I}}_{j,\delta}\left(d,r\right):=d^{\#\mathrm{c}\left(j\right)\delta}{{\mathrm{log}}^{\#\mathrm{a}\left(j\right)\left(1+\delta\right)}r\ },C^{\mathrm{II}}_{j,\delta}\left(d\right):=d^{\#\mathrm{c}\left(j\right)\delta+n\#\mathrm{a}\left(j\right)\left(1+\delta\right)}$
$C^{\mathrm{III}}_{j,\delta}\left(d,r\right):=d^{j\delta}r^{\overline{C}\#\mathrm{a}\left(j\right){\delta}_{m}},C^{\mathrm{IV}}_{j,\delta}\left(d,r\right):=C^{\mathrm{III}}_{j,\delta}\left(d,r\right),A_{j}:=3^{-\#\mathrm{a}\left(j\right)}A$
Property. At the $j$-th interation,
1. The mass of ${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}$ is concentrated over the $O_{j}\in{\mathcal{O}}_{j}$:
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\leq C^{\mathrm{I}}_{j,\delta}\left(d,r\right)\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$ (3.1)
2. The function $f_{O_{j}}$ satisfies
$\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}\leq C^{\mathrm{II}}_{j,\delta}\left(d\right)d^{\#\mathrm{c}\left(j\right)}{\left\|f\right\|}^{2}_{2}$ (3.2)
3. Each $f_{O_{j}}$ satisfies for all $1\leq\rho\leq{\rho}_{j}$.
${\left\|f_{O_{j}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}d^{-\#\mathrm{c}\left(j\right)\left(m-1\right)}{\left\|f\right\|}^{2}_{2}$ (3.3)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C^{\mathrm{III}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ },\ \forall\rho\in\left[1,{\rho}_{j}\right]$ (3.4)
4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
${\left\|f_{O_{j},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{IV}}_{j,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\mathcal{T}}\right\|}^{2}_{2}$ (3.5)
here $\widetilde{f_{O_{j},\ \mathcal{T}}}:={\left(\widetilde{f_{O_{j}}}\right)}_{\mathcal{T}}$ and ${\uparrow}^{j}\mathcal{T}$ denotes the set of wave packets $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ satisfying
${\mathrm{dist}}_{H}\left(\theta,\zeta\right)\leq c_{j}{\rho}^{-1/2}_{j},\ \ {\mathrm{dist}}_{H}\left(T^{L}_{\zeta,w}\left(y_{O_{j}}\right),T^{l}_{\theta,v}\left(y\right)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here
$c_{j}:=\sum^{j}_{i=1}{2^{-\left(j-1\right)/2}C}$
for some absolute constant $C$.
If $j=0$, these properties are trivial. For $j\geq 1$, we will verify these properties by induction on $j$.
Recursion steps. If ${\rho}_{j}\leq r^{{\widetilde{\delta}}_{m-1}}$, then the algorithm terminates. We set $J:=j$ and $\mathcal{O}:={\mathcal{O}}_{j}$. We mark this situation as [tiny]. Otherwise, for each $O_{j}\in{\mathcal{O}}_{j}$ we apply the polynomial partitioning, Theorem 2.29, with degree $d$ to ${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}$. Let ${\mathcal{O}}_{j\mathrm{,cell}}\subseteq{\mathcal{O}}_{j}$ denote the set of all cells for which the cellular case holds and ${\mathcal{O}}_{j\mathrm{,alg}}:={\mathcal{O}}_{j}\backslash{\mathcal{O}}_{j\mathrm{,cell}}$. Then one of the following cases holds:

### v7

@@ 2.7. Transverse Equidistribution Estimates
For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, we define $V^{+}$ by
@@
We cover $B_{\rho}$ by finitely overlapping balls $B$ of radius $r^{1/2+\delta_{m}}$. For a suitable constant $c>0$, if $cB\cap B_{\rho}\cap\mathbf{Z}\neq\emptyset$, we choose a point $x\in cB\cap B_{\rho}\cap\mathbf{Z}$ and denote $V_{B}:=T_{x}\mathbf{Z}$. Let
@@
${\mathbb{T}}_{\bm{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
From the above definition, if $B\in\mathcal{B}_{b}$, then $\angle\!\big(M(\theta),V_{B}^{\perp}\big)$ may be small for some $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$, and consequently the transverse equidistribution estimates may fail for functions concentrated on $\mathbb{T}_{\mathbf{Z},B}$. Fortunately, there are only few sectors $\theta$ for which $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$ for some $v,l$.
@@
$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\bm{\mathrm{Z}},B}\neq\emptyset\right\}$
@@
Suppose $A$ is chosen to be sufficiently large. By the definition of $k$-broad norm and Lemma 2.17,
${\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}=\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Combining this with Lemma 2.11, we have
${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.5)
As shown in [15], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$. The precise statement of the transverse equidistribution estimates for the cone is as follows:
@@
Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
@@
Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$. For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\bm{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
@@
where $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
@@
for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}^{\widetilde{}}$.
@@
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.7)
@@
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.8)
Combining (2.6), (2.7), and (2.8), we obtain
@@
2.8. Polynomial Partitioning
Polynomial partitioning plays a central role in broad estimates. In this subsection, we review the theory of polynomial partitioning. This method is the primary tool used by Guth [5, 6] and Ou–Wang [15] to improve restriction estimates, and it has also led to progress on several related problems, including the Bochner–Riesz conjecture [22] and the Schrödinger maximal estimate [3].
For a polynomial $P:\mathbb{R}^{n}\to\mathbb{R}$, let $\mathrm{cell}(P)$ denote the collection of connected components of $B_{r}\setminus Z(P)$, which partitions the space into finitely many disjoint regions. Each $O\in\mathrm{cell}(P)$ is called a cell cut out by $P$.
@@ 2.8. Polynomial Partitioning
Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$. Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Z}}$, where $\bm{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$. Then, at least one of the following cases holds:
@@
Algebraic case. There exists a $\left(m-1\right)$-dimensional transverse complete intersection $\bm{\mathrm{Y}}\subseteq\bm{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Y}}\right)}$
We now apply polynomial partitioning to the broad estimates. If the cellular case holds, then
$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim d^{m}\|Ef\|^{p}_{BL^{p}_{k,A}(O)}\quad\text{for all }O\in\mathcal{O},$
where $\mathcal{O}$ denotes the collection of cells produced by Theorem 2.20. If the algebraic case holds, then
$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Y}\right)},$
where $\mathbf{Y}$ is an (m1)(m-1)-dimensional transverse complete intersection produced by Theorem 2.20.
By the fundamental theorem of algebra, we deduce that
@@
2.9. Nested Polynomial Wolff Axiom.
To refine the work of Ou–Wang [15], we require the nested polynomial Wolff axioms. We begin by recalling the definitions of grains and multigrains from Section 3 of [12].
Definition 2.22 ([12], Definition 3.3).
A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of radius $r>0$. The dimension of a grain $\left(Z,B\left(y,r\right)\right)$ is that of the variety $Z$, while its degree is that of $Z$ and its scale is $r$.
Definition 2.23 ([12], Definition 3.4).
Let $S=\left(Z,B\left(y,r\right)\right)$ be a grain of dimension $m$. A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belonging to the collection:
Definition 2.24 ([12], Definition 3.5).
A multigrain is a tuple of grains
satisfying
${\mathrm{dim}Z_{i}\ }=i,\quad\forall i$
$B(y_{i},r_{i})\cap B(y_{i+1},r_{i+1})\neq\emptyset,\quad\forall i$
$0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter (ts+1)(t-s+1) is called the multigrain level. The degree of $\overrightarrow{S}$ is defined as $\max_{1\leq i\leq l}Z_{i}$. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=(r_{s},r_{s+1},\dots,r_{t})$. Finally, we write
The restriction conjecture is closely related to a Kakeya-type problem concerning the continuum incidence theory of tubes. More precisely, let $\overrightarrow{S}_{s,t}=(S_{s},\dots,S_{t})$ be a multigrain with $S_{i}=(Z_{i},B(y_{i},r_{i}))$. Suppose $\mathcal{L}$ is a collection of $r_{t}^{-1/2}$-direction-separated rays $L$ satisfying
$L\cap B(y_{i},r_{i})\cap N_{\rho_{j}}Z_{i}\neq\emptyset$
and
$L\cap B(y_{i},2r_{i})\subseteq N_{\rho_{j}}Z_{i}\cap B(y_{i},2r_{i})$
then what is the maximal possible cardinality of $\mathcal{L}$? Understanding this incidence problem is crucial for obtaining effective estimates for $Ef$. It is well known that the restriction conjecture implies the Kakeya conjecture, the latter providing a heuristic bound on the number of incidences between direction-separated tubes.
In [12, 8, 10], a multiscale structure, referred to as nested tubes, is introduced to describe the spatial relationships between tubes at different scales.
Definition 2.25 (Nested Tubes).
and